OpenStudy (anonymous):

the value of tan(2tan^-1 (1/5)-pi/4) is?

OpenStudy (anonymous):

Let\[\theta=2\tan^{-1} (1/5)-(\pi/4)\]Form a triangle with tan theta

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

How what?!

OpenStudy (anonymous):

wrong!

OpenStudy (anonymous):

let x= \[\tan^{-1} \frac{1}{5}\]

OpenStudy (anonymous):

so our expression becomes \[\tan(2x-\frac{\pi}{4})\]

OpenStudy (anonymous):

now apply difference of angle formula

OpenStudy (anonymous):

\[\tan(A-B) = \frac{\tan(A)-\tan(B) } {1+\tan(A)\tan(B) } \]

OpenStudy (anonymous):

so we get

OpenStudy (anonymous):

\[\frac{\tan(2x) -1}{1+\tan(2x)} \]

OpenStudy (anonymous):

but ,\[\tan(2x) = \frac{\sin(2x)}{\cos(2x)} = \frac{2\sin(x)\cos(x)}{\cos^2(x)-\sin^2(x)} \]

OpenStudy (anonymous):

now, we go back to our substitution x= tan^-1(1/5) take tan of both sides , and we get tan(x) = 1/5 now we can draw up a general right angle triangle and mark an angle x , and fill in the lengths 1 and 5 for thre opposite and adjacent sides respectively remember tan = opposite/adjacent

OpenStudy (anonymous):

note the hypotenuse is sqrt(26)

OpenStudy (anonymous):

now from that triangle we can find the value of sin(x) and cos(x) , the then you just plug them into the expression we had above

OpenStudy (anonymous):

so \[\sin(x) = \frac{1}{\sqrt{26}}\] \[\cos(x) = \frac{5}{\sqrt{26}}\] just by using the definitions

OpenStudy (anonymous):

so , our whole expression was\[\tan(2x- \frac{\pi}{4}) = \frac{ \frac{ 2\sin(x)\cos(x) }{\cos^2(x)-\sin^2(x) } -1}{ 1+ \frac{2\sin(x)\cos(x)}{\cos^2(x)-\sin^2(x)}}\]

OpenStudy (anonymous):

then sub in the values for sin(x) and cos(x) above, and simplify

OpenStudy (anonymous):

you can do that

OpenStudy (anonymous):

good. it worked out well. thanx man!